OpenStudy (anonymous):

hey dose anyone know how to do the problem 2y^3 - 128 the formula for this that I am trying to use is difference of squares A^3 - B^3 = (A-B) (A^2 + AB + B^2) the direct5ions fo this problem are to factor out completely i am very confused please help!

OpenStudy (loser66):

first off, factor 2 out, you have 2(y^3-64) , ok now, right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

good!

OpenStudy (anonymous):

okay so whats the next step and your getting the 64 from 128 right?

OpenStudy (anonymous):

could you use the draw thing if you could show your steps that would help a lot.

jimthompson5910 (jim_thompson5910):

hint: 64 = 4^3 so 2(y^3-64) = 2(y^3-4^3)

OpenStudy (loser66):

:) @huntergirl1 When you said yes, I thought you got all of them since you gave me the formula. You just apply the formula you post to get the answer.

OpenStudy (anonymous):

okay so i end up with what i started with which is great because that was the correct answer and yes sorry i was having a dyslexia moment with the 64 thing i understand what you got and this is what i got... 2(y-4) (y^2 +4y +16) 2(y^3 + 4y^2 +16y -4y^2 -16y-64) i simplied with like terms and then i got this. 2(y^3-64)= 2y^3-128 the original problem

OpenStudy (anonymous):

unless i am completely wrong

jimthompson5910 (jim_thompson5910):

2y^3 - 128 2(y^3 - 64) 2(y^3 - 4^3) 2(y-4)(y^2 + y*4 + 4^2) ... apply the difference of cubes rule here 2(y-4)(y^2 + 4y + 16) and that's as far as you can go

OpenStudy (anonymous):

so yes this is correct then?

jimthompson5910 (jim_thompson5910):

the steps after 2(y-4)(y^2 + 4y + 16) appear to be you expanding everything out again, which is the opposite of what they want. They want you to factor.

OpenStudy (anonymous):

okay then so all i need is the set up part?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

so im leaving it at 2(y-4) (y^2 +4y +16)

OpenStudy (anonymous):

or is that not fully factored out?

jimthompson5910 (jim_thompson5910):

yeah that's the fully factored form

OpenStudy (anonymous):

okay thanks!! (: and also would you be up to helping me with one more? rs^2 +64? same directions and would i start by factoring the r out completely?

jimthompson5910 (jim_thompson5910):

I don't see an 'r' in the second term

jimthompson5910 (jim_thompson5910):

is there one?

OpenStudy (anonymous):

sorry typo that 64 is supposed to have an r on it as well |

OpenStudy (anonymous):

rs^3 + 64r should i take and factor out the r first?

jimthompson5910 (jim_thompson5910):

yes, and what do you get when you do so?

OpenStudy (anonymous):

r(s^3 +4^3)

OpenStudy (anonymous):

is that on the right track?

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

then you'll use the sum of cubes factoring formula

OpenStudy (anonymous):

and once it is plugged into the formula then i am done right? with the factoring?

jimthompson5910 (jim_thompson5910):

yes, tell me what you get

OpenStudy (anonymous):

okay i got r(s+4) (s^2- 4s+ 16)

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

okay and then i went on and i did this r(s^3 - 4s^2 + 16s+ 4s^2 -16s + 64) I'm going to combine like terms and then end up with... r(s^3 + 64) right or did i go to far

jimthompson5910 (jim_thompson5910):

why are you expanding again? they just want you to factor

jimthompson5910 (jim_thompson5910):

r(s+4) (s^2- 4s+ 16) is the fully factored form when you got to r(s^3 - 4s^2 + 16s+ 4s^2 -16s + 64), it's correct but not what they want. They don't want you to expand. I'm assuming they simply said "factor the expression"

OpenStudy (anonymous):

they just said fator ompletely but okay so then leave it at that

jimthompson5910 (jim_thompson5910):

yeah, r(s^3 - 4s^2 + 16s+ 4s^2 -16s + 64) takes a step backward from factoring so leave it as r(s+4) (s^2- 4s+ 16)

OpenStudy (anonymous):

okay then but if i was going to check this because i have a quiz on this information tomorrow the I would continue to expand? i would show my answer as-->r(s+4) (s^2- 4s+ 16) then expand to check my answr to make sure that it matches the original problem right? .

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

okay cool!! thank you soooo much!!! i will not torture you anymore with my questions thanks for the help!!

jimthompson5910 (jim_thompson5910):

lol you're ok, no worries. I'm glad I could help out